OpenStudy (anonymous):

-(tan^2 -sin^2 -cos^2-sec^2)=

OpenStudy (anonymous):

\[-\tan^2x + \sin^2x + \cos^2x +\sec^2x\] \[\sin^2x + \cos^2x +\sec^2x - \tan^2x\] Remember: \[\sec^2x - tanx = 1\] \[\sin^2x + \cos^2x = 1\] Use these formulae and get the answer...

OpenStudy (anonymous):

so answer would be 2

OpenStudy (anonymous):

You are right..

OpenStudy (anonymous):

There it will come tan^2x in place of tanx...